OpenStudy (anonymous):

Write in exponential form log^9 729=3

OpenStudy (anonymous):

\[\left( \log _{9} 279 \right) =3\] Put a base of 9 for each side \[9^{(\log _{9} 729)} = 9^{3}\] The 9 and the base 9 in the log cancel out so you're left with \[9^{3}=729\] That is the Exponential form